# Enabling Delayed-Full Charging Through Transformer-Based Real-Time-to-Departure Modeling for EV Battery Longevity

### v1

@@ Conclusion
This paper presents a Transformer-based TTD modeling framework as the foundation for DFC, a strategy aimed at reducing battery degradation and extending EV battery life. Unlike prior work that relies on historical patterns, our approach uses real-time contextual signals from unobtrusive smartphone sensing and models departure as a survival process on discretized time grids. In a large-scale in-the-wild study with 93 participants and 3,906 daily sequences, our method outperforms historical and context-aware baselines, reducing MAE by 14.4%. Ablation studies highlight the role of design considerations in robust temporal modeling, while fine-tuning enables further personalization. The proposed dataset and framework establish a benchmark for real-time departure prediction in sustainable EV smart charging, supporting both battery longevity and carbon neutrality.

### v2

@@ Conclusion
This paper presents a Transformer-based TTD modeling framework as the foundation for DFC, a strategy aimed at reducing battery degradation and extending EV battery lifetime. Unlike prior work that relies on historical patterns, our approach uses real-time contextual signals from unobtrusive smartphone sensing and models departure as a survival process on discretized time grids. In a large-scale in-the-wild study with 93 participants and 3,906 daily sequences, our method outperforms historical and context-aware baselines, reducing MAE by 14.4%. Ablation studies highlight the role of design considerations in robust temporal modeling, while fine-tuning enables further personalization. The proposed dataset and framework establish a benchmark for real-time departure prediction in sustainable EV smart charging, supporting both battery longevity and carbon neutrality.